BARTLETT’S FORMULAE—CLOSED FORMS AND RECURRENT EQUATIONS

GEORGI N. BOSHNAKOV

Institute of Mathematics, Acad. G.Bonchev str., 8, 1113 Sofia, Bulgaria

(Received November 21, 1994; revised May 8, 1995)

Abstract. We show that the entries of the asymptotic covariance matrix of the sample autocovariances and autocorrelations of a stationary process can be expressed in terms of the square of its spectral density. This leads to closed form expressions and fast computational algorithms.

Key words and phrases: Bartlett’s formula, ARMA, sample autocovariances, sample autocorrelations.